### v1

# FLEX: A Large-Scale Multi-Modal Multi-Action Dataset for Fitness Action Quality Assessment

Abstract
With the increasing awareness of health and the growing desire for aesthetic physique, fitness has become a prevailing trend. However, the potential risks associated with fitness training, especially with weight-loaded fitness actions, cannot be overlooked. Action Quality Assessment (AQA), a technology that quantifies the quality of human action and provides feedback, holds the potential to assist fitness enthusiasts of varying skill levels in achieving better training outcomes. Nevertheless, current AQA methodologies and datasets are limited to single-view competitive sports scenarios and RGB modality and lack professional assessment and guidance of fitness actions. To address this gap, we propose the FLEX dataset, the first multi-modal, multi-action, large-scale dataset that incorporates surface electromyography (sEMG) signals into AQA. FLEX utilizes high-precision MoCap to collect 20 different weight-loaded actions performed by 38 subjects across 3 different skill levels for 10 repetitions each, containing 5 different views of the RGB video, 3D pose, sEMG, and physiological information. Additionally, FLEX incorporates knowledge graphs into AQA, constructing annotation rules in the form of penalty functions that map weight-loaded actions, action keysteps, error types, and feedback. We conducted various baseline methodologies on FLEX, demonstrating that multimodal data, multiview data, and fine-grained annotations significantly enhance model performance. FLEX not only advances AQA methodologies and datasets towards multi-modal and multi-action scenarios but also fosters the integration of artificial intelligence within the fitness domain. Dataset and code are available at here.
1 Introduction
Physical exercise—particularly weight training—serves multiple purposes, ranging from supporting healthy cognitive function to preserving musculature and bone density. However, weight training also carries a risk of physical injury, often resulting from improper form, posture, or movement. Therefore, proper coaching and maintaining correct workout form are crucial to prevent injuries and fully reap the benefits of resistance training.
Figure 1: An overview of the FLEX dataset. FLEX dataset consists of a core group of 38 subjects, each performing 20 different fitness actions, repeating each action 10 times. Each action repeat was recorded from 5 viewpoints, & sEMG signals and physiological parameters (heart rate, breath rate) were simultaneously collected along with videos. The data annotations contain rich text information such as action knots (AK), error types (ET), & action feedback. (Zoom in for the best view.)
Prior research has demonstrated that computer vision and AI-based solutions can assist in this area. Specifically, videos of individuals performing exercises can be analyzed by models to detect errors in workout form. In the video understanding literature, this task is commonly referred to as Action Quality Assessment (AQA)—the process of quantifying how well an action, or in this context, an exercise, is performed. To train AQA models, datasets typically consist of videos of people performing actions along with corresponding annotations such as action quality scores or error labels.
However, existing fitness datasets are limited in several key aspects: 1) they cover a narrow range of exercise actions; 2) often exclude weight-loaded exercises; and 3) lack diversity in subject demographics and skill levels. To address these gaps, we introduce the FLEX dataset—a high-quality resource for exercise assessment and coaching, designed with the following desirable characteristics:
  • Multiview: High-precision markerless MoCap system used to capture video data from 5 viewpoints
  • Multimodal: A Custom wearable vest with a surface EMG system was used to record muscle activation and physiological signals synchronized with video and 3D joint data. A 3D scanner and body composition analyzer captured each subject’s point cloud and body metrics.
  • Largescale: 40+ hours of high-quality multimodal dataset—20 weight-loaded fitness actions, each performed 10 times by 38 subjects. With 7,512 samples, it is the largest dataset in AQA.
  • Weight loading: FLEX includes 20 common weight training exercises to help prevent injuries.
  • Fitness Knowledge Graph (KG): We defined key action steps and built a fitness-action knowledge graph linking action keysteps, error types, and feedback.
  • Multirepetition: Each action is performed 10 times, enabling applications like AQA-based early failure prediction in fitness and sport.
  • Subject skills-level diversity: To enable assessment of a diverse range of skill levels, we recruit 38 human subjects from 3 skill levels, namely: Novice, Amateur, and Expert.
  • Expert annotations and Quality Checks: Guided by our Fitness Knowledge Graph, professionals meticulously annotated the dataset with action keysteps, error types, AQA scores, and corrective feedback. Multilevel verification and quality checks ensured data accuracy.
FLEX dataset, with these comprehensive attributes, represents a high-quality resource that supports the development of the next generation of sophisticated AI models for fine-grained action assessment and coaching and unlocks new capabilities like Video2EMG estimation for precise muscle-level intervention and AQA-based early failure prediction in fitness and sports.
2 Related Work
Action Quality Assessment Datasets. Datasets are essential to machine learning and AQA, supporting model training, evaluation, and deployment. Since the foundational work in 2014, many high-quality AQA datasets have emerged across diverse domains, including sports pirsiavash2014assessing ; parmar2017learning , healthcaregao2014jhu ; vakanski2018data ; capecci2019kimore , fitnessparmar2022domain ; li2024egoexo ; ogata2019temporal , industrial trainingsener2022assembly101 , and generative AI evaluationchen2024gaia . Early datasets, such as MIT-Dive, MIT-Skate, UNLV-Dive, and UNLV-Vault contained only single-action categories. This was followed by multi-action and multimodal datasets such as AQA-7 and MTL-AQA (contains action scores, finegrained action class labels, verbal description of action quality).
Datasets like FineDiving split the actions into clips, representing a trend toward more detailed annotation. Furthermore, the emergence of datasets like Egoexo-Fitness and LucidAction represents a shift from crawling online videos toward in-person collection of AQA data, despite being resource-intensive. This approach guarantees control over both the dataset scale and quality. Advances in AQA datasets have contributed to developing AQA methodologies, with multimodal, fine-grained, and multi-action datasets allowing models to be more comprehensive and fine-grained, improving model generalization capabilities. However, existing datasets predominantly focus on competitive sports actions, with only a few dedicated to the fitness domain (for details see Table 1). Moreover, current fitness datasets primarily include self-loaded actions ogata2019temporal ; li2024egoexo , exhibit narrow action categories parmar2022domain , or are constrained to RGB modalities ogata2019temporal ; parmar2022domain ; li2024egoexo , thereby limiting the application and innovation.
Table 1: Comparison of FLEX dataset with representative AQA datasets in terms of sample size, action types, views, modalities, annotations, & sources. V: Video, Sk: Skeleton, sE: sEMG, P: Physiological Info., S: Score, G: Grade, A: Action, F: Formation, D: Description, Fb: Feedback.
Dataset Year Sample Type View Modality Annotation Domain Source
MIT-Divepirsiavash2014assessing 2014 159 1 1 V S Sport Web
MIT-Skatepirsiavash2014assessing 2014 150 1 1 V S Sport Web
UNLV-Diveparmar2017learning 2017 370 1 1 V S Sport Web
UNLV-Vaultparmar2017learning 2017 176 1 1 V S Sport Web
AQA-7parmar2019action 2019 1189 7 1 V S, A Sport Web
MTL-AQAparmar2019and 2019 1412 2 1 V S, A, D Sport Web
Waseda-Squatogata2019temporal 2019 2001 1 1 V, Sk E Fitness Camera
Fis-Vxu2019learning 2019 500 1 1 V S Sport Web
TASD-2gao2020asymmetric 2020 606 2 1 V S, A Sport Web
Rhy.Gym.zeng2020hybrid 2020 1000 4 1 V S, A Sport Web
FR-FSwang2021tsa 2021 417 1 1 V G, A Sport Web
Fitness-AQAparmar2022domain 2022 13049 3 1 V G, A Fitness Web
FineDivingxu2022finediving 2022 3000 52 1 V S, A Sport Web
LOGOzhang2023logo 2023 200 12 1 V S, A, F Sport Web
FineFSji2023localization 2023 1167 1 1 V S, A Sport Web
PaSkgao2023automatic 2023 1018 1 1 V S Sport Web
EgoExo-Fitnessli2024egoexo 2024 6131 12 6 V S, A, D Fitness Camera
LucidActiondong2024lucidaction 2024 6702 8 8 V, Sk S, A Sport MoCap
Ours FLEX 2025 7512 20 5 V, Sk, sE, P S, A, D, E, Fb Fitness MoCap
To bridge those gaps and stimulate innovation in AI-based Fitness coaching, we develop FLEX—first AQA dataset that provides multi-view videos with synchronized 3D pose, sEMG, and physiological information. Furthermore, we develop detailed, professional, comprehensive annotation rules for fitness actions that integrate multiple knowledge sources. These rules form a KG containing Action Keysteps, Error Types, and Action Advice, making the annotation rich in paired textual information.
Electromyography (EMG) Datasets. EMG records electrical activity from skeletal muscles to assess their activation levels, motor coordination, and detect abnormalities. EMG devices are broadly classified into surface EMG (sEMG), which uses non-invasive skin electrodes, and needle EMG (nEMG), which involves inserting fine needles into muscle tissue alcan2023current . Owing to its non-invasive nature, existing EMG datasets predominantly utilize sEMG for applications in biomechanical analysiszhang2017extracting ; hug2019individuals ; mileti2020muscle ; jarque2020large ; santuz2022sex ; khan2020semg ; dimitrov2023hig ; hamard2023comparison ; wang2023wearable , hand gesture estimationatzori2014electromyography ; khan2020semg ; jarque2020large ; liu2021neuropose ; simpetru2022sensing ; ozdemir2022dataset ; salter2024emg2pose , etc.
Existing EMG datasets mainly link signals to motion categories, overlooking their connection to action quality. Since EMG reflects neuromuscular control and muscle function—key in fitness—activation patterns serve as objective quality markers. To address this, FLEX is the first AQA dataset to integrate EMG, using a 16-channel sEMG system fastmove at 2,000 Hz to record target muscle activity during fitness movements. sEMG and MoCap were synchronized for precise temporal alignment.
3 FLEX Dataset
We present FLEX, a multimodal, multiview dataset that addresses key limitations in existing fitness AQA datasets: limited exercise variety, lack of weight-loaded actions, and low skill diversity. To our knowledge, FLEX is the first weight training dataset with multimodal data, multiview capture, detailed exercise procedures, standardized assessment rules, and corrective feedback. This section covers data collection, annotation, quality checks, procedures, and evaluation standards.
3.1 Preliminary Setup
Figure 2: Comparison between FLEX and Egoexo-fitness datasets. EWL: equipment (barbell, dumbbell, etc.)-based weight loading; IR: level of risk of injury.
Dataset Exercises EWL IR
Fitness-AQAparmar2022domain 3 high
EgoExoli2024egoexo 12 low
Ours FLEX 20 high
Action/Exercise Selection.
Bodyweight exercises require no equipment, making them simpler and less injury-prone. In contrast, weight-loaded exercises involve lifting external weights, targeting specific muscles, and following structured progression, but carry a higher injury risk if done improperly. Thus, proper technique and posture are crucial in weight training to prevent injuries. To support this, we carefully selected 20 common exercises based on the following criteria: 1) Include exercises targeting both upper and lower body muscle groups. 2) Cover complex, injury-prone joints, such as the shoulders. 3) Feature a variety of higher-risk free-weight exercises. 4) Focus on widely practiced exercises to ensure broad relevance, aligning with commonly recommended training regimens. A comparison of our dataset with Fitness-AQA and EgoExo-Fitness is provided in Figure 2.
Figure 3: Comparison of ability level diversity of the subjects in AQA datasets. N: Novice, A: Amateur, P: Professional.
Dataset N A P
Sport Domain Datasetspirsiavash2014assessing ; parmar2017learning ; xu2022finediving
Waseda-Squatogata2019temporal
Fitness-AQAparmar2022domain
EgoExo-Fitnessli2024egoexo
LucidaActiondong2024lucidaction
Ours FLEX
Diverse Subject Pool. To capture diverse skill levels, we recruited subjects across Novice, Amateur, and Expert skill brackets through surveys at our institutions and local gyms, receiving over 60 registrations from students and coaches. Each applicant completed an online questionnaire, an offline ability test, and 3D/body composition analysis to assess weight-bearing capacity and skill level. We selected 38 male subjects: 10 experts (ID: P01–P10), 8 amateurs (ID: A01–A10, excluding A06 and A09), and 20 novices (ID: N01–N10). Unlike existing datasets, ours spans all skill levels, as shown in Figure 3.
Recording Multiple Continuous Exercise Repetitions. To support modeling the temporal evolution of action quality over time, each subject was required to perform 10 repetitions per action. In comparison, existing datasets lack this multirepetition capture of weight training.
Human subject safety. To ensure the safety of human subjects during data collection of risk-prone exercises, they were asked to perform exercises at 80% of their maximum weight-loaded capacity. Additionally, all subjects were covered by sports injury insurance, with on-site safety supervisors and emergency medical kits available throughout the process. Our data collection process was reviewed and approved by the Institutional Review Board (IRB) of Beijing Sport University (Approval No.: 2023313H). All participants provided written informed consent prior to participation.
3.2 Data Collection
Multiview Capture.
Fitness actions involve people in convoluted positions and postures (not captured in typical pose estimation datasets). Additionally, in weight training exercises, essential body parts and joints might be occluded if the scene is captured from any single camera angle. To obtain more complete (although not fully complete) groundtruth motion capture, we record the human from multiple camera angles using a high-precision markerless motion capture systemfastmove_gait capable of acquiring multi-view videos and 3D skeletal points. This system comprises four synchronized ZCAM E2M4 cinema cameraszcam , each equipped with an Olympus 14–150 mm lens (fixed at 14 mm)lens , recording at 4K resolution (3840$\times$2160 pixels) with 120 FPS. To replicate recording using a typical smartphone and improving the accessibility, we additionally incorporated a smartphone view captured with a OnePlus 7phone at 1080P resolution (1920$\times$1080 pixels) and 60 FPS.
Figure 4: Data collection environment. Four cinema cameras and one smartphone were fixed at the four corners of the collection area. Video, sEMG, heart rate, and breath rate are recorded synchronously during collection.
Multimodal information. Existing datasets predominantly include modalities such as imagespirsiavash2014assessing , textsparmar2019and , skeletal pointscapecci2019kimore , and audioparmar2021piano , with limited exploration of other potentially valuable physiological modalities. Physiological modalities, especially EMG, which directly reflects deep muscular activity and is naturally related to movement or action quality, provide particularly relevant and deep feedback for AQA in fitness scenarios. Thus, in the FLEX dataset, we record: 1) EMG signals using a 16-channel FastMove sEMGfastmove device operating at a sampling rate of 2000 Hz. sEMG captures muscle activity. 2) Heart rate, collected with the help of a customized wearable vest, provides insight into a person’s exertion rate. 3) Respiratory rate: Exercise requires proper and procedural breathing. Our exercise procedure base and rules include breathing as an AQA criterion.
Motion capture environment and sensor placements. All devices were installed within a controlled space measuring 5m×$\times$5m×$\times$2m (as shown in Figure 4). The four cinema cameras for motion capture were fixed at each corner of the square space, while the smartphone was positioned at the main viewpoint. All four sides of the controlled space were covered with curtains to reduce environmental interference further. The placement of the sEMG sensors was adjusted based on the specific target muscles involved in each action. Each subject wore the customized vest tightly to ensure accuracy in data acquisition.
Collection process. To begin the data collection process, each subject was first informed of the actions to perform. Subsequently, a data collection staff assisted human subjects in wearing the customized vest and attaching the sEMG sensors to the target muscle regions. Notably, in contrast to data collection strategies adopted by the FLAG3Dtang2023flag3d and Egoexo-Fitnessli2024egoexo datasets, subjects in our study were required to enter the collection area individually and sequentially—to avoid observing others or receiving detailed instructions from the data collection staff. This approach prevented external factors from influencing the subjects’ performance levels. Additionally, no textual action descriptions were provided during the data collection; instead, a concise demonstration was shown briefly before collecting started, ensuring that subjects performed the actions based solely on their own experience. After the sensor attachment, the subject entered the collection area, facing the curtain between Viewpoints 1 and 2. Upon receiving the staff’s signal to begin the test, the subject independently performed the specified action 10 times, after which the staff recorded the corresponding weight data. The whole process was overseen by the authors.
Rest periods in between data collection. Considering the difficulty of performing 20 consecutive actions, particularly for the 20 novice subjects, only professional-level subjects performed all 20 consecutively. The remaining 28 subjects completed their collections separately over multiple days to ensure high-quality data, grouped by targeted muscle regions.
3.3 Data Annotation
The value of a high-quality dataset relies not only on the accuracy and completeness of the raw data but also significantly on the precision and consistency of the annotations. During the data collection phase, we employed multiple devices to acquire multimodal raw data across diverse actions, maintaining close monitoring throughout to guarantee data quality. As elaborated in the following sections, the FLEX dataset implemented stringent measures in formulating annotation rules and selecting annotators to ensure further rigorous annotation.
3.3.1 Developing Standardized Fitness Rules and Knowledge-base
Unlike competitive sports (e.g., diving, figure skating, gymnastics) with clearly defined action objectives and scoring criteria, fitness actions aim to serve diverse goals such as muscle growth, fat loss, and rehabilitation. Consequently, different individuals might adopt varying evaluation standards for the same action. To guarantee high-quality annotation for the FLEX dataset, we extensively gathered existing fitness action standards from multiple sources, including national standards, fitness associations, professional literature, fitness applications, and influential fitness content creators. We established a precise and objective annotation system through comprehensive integration and refinement.
Sources of Fitness Action Standards. Specifically, our annotation rules primarily drew upon the "National Occupational Skill Standards—Social Sports Instructor (Occupation Code: 4-13-04-01)" jointly formulated by the Ministry of Human Resources and Social Security of the People’s Republic of China and General Administration of Sport of Chinastandard2020 . Additionally, we referenced the "Occupational Competency Training Textbook for Social Sports Instructors—Fitness Coaches (with Technical Action Videos)" published by the Human Resources Development Center of the General Administration of Sport of ChinaHRDSport_Toolkit , the "Fitness and Bodybuilding Tutorial" from Beijing Sport Universitybsu2013 , and "Joe Weider’s Bodybuilding System"joe .
Developing Standardized Annotation Rules through Multi-Source Integration and Expert Review. We thoroughly reviewed all sources, meticulously documenting essential information relevant to the collected actions, emphasizing target muscles, action descriptions, and potential error types. Through rigorous comparison and integration, we distilled these insights into annotation rules that adhere to existing standards while accommodating practical training scenarios. These rules underwent rigorous verification by experts from Beijing Sport University, ensuring both theoretical rigor and enhanced consistency and practicality in annotations.
Action Segmentation and Keystep Modeling. We segmented each collected action into three distinct phases based on major limb trends: Preparation, Concentration, and Eccentric. Within each phase, the main limb actions were further subdivided into several action key-steps (AK), with integrated descriptions assigned accordingly. Combining these AKs systematically enabled the generation of textual descriptions for diverse fitness actions.
Error Type Identification and Feedback Design. Additionally, we identified specific error types (ET) associated with localized joint actions within each major limb action and provided corresponding corrective feedback suggestions.
Knowledge Graph Construction. Consequently, we constructed a fitness action knowledge graph (see in the Appendix) comprising actions, action keysteps (AK), error types (ET), feedback, and their interrelationships.
Scoring System for Annotation Efficiency. Furthermore, different error types were assigned distinct weights within a penalty-based scoring system, enabling annotators to efficiently identify and label errors from action clips. The final scores were obtained through cumulative penalties, ensuring both efficiency and consistency throughout the annotation process (provided in the Appendix).
Figure 5: Annotation Process. The annotators we recruited were trained according to the sources of the annotation guidelines and underwent centralized training to ensure they thoroughly understood the rules. The video data was segmented following predetermined criteria, and a two-stage annotation process was implemented to reduce annotation errors and mitigate subjective bias.
3.3.2 Annotation Process
We followed the steps for annotating our dataset.
  1. 1.
    Annotator Recruitment. We carefully recruited 16 fitness professionals and practitioners to annotate our dataset. For details on the recruitment procedure, please refer to the Appendix.
  2. 2.
    Action Segmentation. During the data segmentation phase, annotators segmented videos containing repeated actions into individual action samples according to the previous standards and guidelines and verified each sample.
  3. 3.
    Task Organization and Annotator Training. In the formal annotation phase, we organized the tasks into several work periods based on the target muscle groups of the actions and provided centralized training for 16 annotators at the beginning of each period to ensure a thorough understanding of the annotation guidelines.
  4. 4.
    Annotator Grouping by Experience. Based on their fitness experience, the annotators were divided into two groups: the first group—comprising trainers with less than three years of experience and master’s students—handled the initial annotation (Stage 1), while the second group—consisting of trainers with over three years of experience and doctoral students—was responsible for annotation verification (Stage 2).
  5. 5.
    Two-Stage Annotation & Verification Workflow. In Stage 1, each sample was initially annotated for error types by three different annotators from the first group using both video and 3D pose information. In Stage 2, these three sets of initial annotations were reviewed by three annotators from the second group. If any issues were identified, the final annotation was determined by a majority vote among these three reviewers to mitigate subjective bias.
  6. 6.
    Quality Control Measures. Authors periodically monitored the quality of the annotated data to ensure that the dataset maintained a high degree of reliability and academic value for subsequent analyses.
Dataset Statistics and Visualization. In total, we collect 7512 samples from 38 subjects doing 20 exercises, each for 10 repetitions. We have provided further dataset stats along with WordClouds and visualized samples in the Appendix.
3.4 FLEX VideoQA Dataset
We leverage LLM to generate Questions and Answers based on our Fitness Knowledge Graph and annotations related to error types, action description and feedback. These QA sets are then manually verified and corrected if needed. More details are provided in the Appendix.
4 Experiments
FLEX dataset’s Multimodal data, Fitness Action Knowledge Graph, and the fine-grained annotations not only support detailed action quality assessment but also serve as a rich resource for diverse research scenarios such as chatbot-based fitness assessment and coaching (involves video question answering) and crossmodal signal estimation. To foster future research on FLEX, we have designed and conducted multiple benchmark evaluations on the dataset to demonstrate its broad application potential.
4.1 Action Quality Assessment
Baselines. In this section, we develop the following baselines by extending the CoRe model yu2021group .
  1. 1.
    Multiview (MV): This model uses multi-view video, based on the hypothesis that combining perspectives provides a more complete view of posture and movement, enhancing AQA performance.
  2. 2.
    Multiview incorporating explicit Pose information (MV+Pose): We hypothesize that pose features can enhance the video-based AQA model. Thus this model builds on MV model and incorporates pose features extracted via an ST-GCNyan2018spatial pre-trained on NTU-RGB+Dshahroudy2016ntu .
  3. 3.
    Multiview incorporating explicit Pose information and Muscle-level Physiological information (MV+Pose+EMG): sEMG offers direct insight into muscle activity, valuable for AQA. This model extends MV+Pose by integrating sEMG. We do this by computing relative muscle contribution and concatenating it with video and pose features.
Implementation details. Provided in the Appendix. Full Codebase has been released.
Metrics. Following standard practice, we report Spearman’s rank correlation ($\rho$) and $R-l2$ metrics.
Figure 6: Performance of AQA models on FLEX dataset. $R-l2(\times 100)$
AQA Model $\rho\uparrow$ $R-l2\downarrow$
Single viewyu2021group 0.8069 1.7582
Multiview (MV) 0.8974 0.9095
MV + Pose 0.9003 0.8920
MV + Pose + EMG 0.9019 0.8877
Results. For a fair comparison, we keep the architecture, loss function and training strategy same across models. Results are summarized in Figure 6. On our FLEX dataset, the single view model achieves an $\rho$ of 0.8069 and an $R\mbox{-}l2$ of 1.7582, demonstrating the efficacy of contrastive regression modeling. When multiview data is leveraged—MV model—the model attains substantial gains on both $\rho$ rises to 0.8974 and $R\mbox{-}l2$ falls to 0.9095—indicating that the second viewpoint enriches appearance information in the action space and enables more robust similarity discrimination. Further integrating skeletal point information (MV+Pose) boosts performance further ($\rho$ = 0.9003, $R\mbox{-}l2$ = 0.8920), underscoring the complementary value of high-level structural motion representations for capturing movement trends and amplitude. Finally, by introducing sEMG features (MV+Pose+EMG), the model achieves its best results, $\rho$ reaches 0.9019, and $R\mbox{-}l2$ drops to 0.8877. These findings suggest that the muscle-level control signals provided by sEMG can supplement external motion trajectories by capturing internal physiological states, thereby furnishing a more mechanistic foundation for action quality assessment.
4.2 Action Quality Understanding
We envision a future where people can use chatbots as AI-Coaches, who can understand and monitor peoples’ exercise posture and movements and provide corrective feedback on it. Vision-Language Models (VLMs) or Multimodal Large Language Models (MLLMs) are a natural choice to implement such AI-Coach because their capability to conduct multimodal communication and processing. However, out of the box, these models might not have specialized skills/knowledge to perform fitness action quality assessment. We believe our multimodal data, the fitness action knowledge graph, and the fine-grained annotations in the FLEX dataset can bridge this gap. To that end, we create FLEX-VideoQA dataset (subsection 3.4) containing fitness-based video question-answering samples. This dataset unlocks the integration of VLMs/MLLMs in AQA research.
Figure 7: Performance of Qwen2.5-VL 3B on FLEX-VideoQA dataset. BF1: BERTScore F1, SFT: Supervised Fine-Tuning. $R-l2(\times 100)$
Metrics Pretrained Prompt SFT
BLEU $\uparrow$ 0.0388 0.0415 0.1970
ROUGE-L $\uparrow$ 0.2012 0.2128 0.4010
METEOR $\uparrow$ 0.3049 0.3206 0.4688
CHRF++ $\uparrow$ 36.1340 39.5308 52.0005
BF1 $\uparrow$ 0.1664 0.1666 0.4390
$R-l2\downarrow$ 3.1183 3.0942 3.0592
Setup. Noting its strong performance, we choose Qwen2.5-VL-3Bbai2025qwen2 , a recently released multimodal chatbot model to develop our AI-driven Fitness Coach. Note that other VLMs can also be used. We evaluated its performance under three configurations: 1) Off-the-Shelf mode: Using the original pretrained weights. 2) Rules-based Prompting mode: Providing standardized action guidelines during the evaluation phase. 3) Supervised fine-tuning mode: Applying a LoRA strategy to fine-tune the entire model on top of the pretrained weights. Further implementation details are provided in the Appendix.
Metrics. To comprehensively assess Qwen’s outputs, we employed BLEU, ROUGE-L, METEOR, CHRF++, and BERTScore-F1, which are commonly used in LLM research, as metrics of generated text, alongside $R\mbox{-}l2$ to evaluate prediction-score accuracy.
Results. Analysis of the results in Figure 7 shows that supervised fine-tuning delivers an order-of-magnitude improvement in text-generation quality compared to the pretrained model, yet only yields marginal improvements in consistency of action scoring. The results show significant improvement in semantic alignment between fitness terminology and visual entities—not covered in mainstream video understanding datasets, but covered in our dataset. However, $R\mbox{-}l2$ decreases by less than 2%, indicating that the current fine-tuning strategy optimizes the language-generation head and contributes minimally to the regression head, which relies on fine-grained pose differences. While current VLMs still underperform specialized AQA models in precise action scoring, their ability to generate natural, professional action descriptions offers a viable path for integrating VideoQA task with AQA task.
4.3 Novel Task–Video2EMG
sEMG provides feedback on muscle functional state and correlates with action quality. We envision a future where EMG signal could be estimated directly videos. Such technology would enable provide precise—muscle-level—feedback and intervention. To that end, we introduce a new task of estimating sEMG signals from videos. Our FLEX dataset with paired recordings of videos and EMG signals can enable this new research direction.
Video2EMG Baseline Model. Inspired by the emg2posesalter2024emg2pose framework, we designed an end-to-end visual regression model combining a ResNet backbone with an LSTM.
Implementation details. Provided in the Appendix. Code for this model has been released.
Metrics. To evaluate the Video2EMG model’s regression performance, we report mean absolute error (MAE) and root mean square error (RMSE), capturing average error and its variability, respectively.
Figure 8: Performance of Video2EMG models on FLEX dataset.
Video2EMG Model MAE $\downarrow$ RMSE $\downarrow$
ResNet-LSTM 0.1655 0.2133
Results. After 150 training epochs, the Video2EMG model exhibited strong regression performance (see in Figure 8). Following inverse normalization, it achieved an MAE of 0.1655 and an RMSE of 0.2134, demonstrating accurate sEMG value prediction across most samples while keeping error variability within a reasonable range. These results serve as baseline for future work.
5 Limitations and Future Work
Limitations. Despite our efforts to broaden recruitment, several contextual factors limited the inclusion of female participants in the FLEX dataset. The institute’s remote location, local gender imbalance, and sociocultural norms posed significant challenges. As a result, the dataset comprises 38 male participants. We acknowledge this limitation and are committed to improving gender diversity in future studies through expanded recruitment strategies and partnerships.
Future Work. FLEX dataset provides the computer vision community with novel, high-quality multimodal resources and establishes an extensible paradigm for future fitness datasets. Our preliminary VLM and Video2EMG experiments further demonstrate the promise of cross-modal research and integrating AQA models with vision–language models to realize a unified “action agent.”
6 Conclusion
In this paper, we propose FLEX, a large-scale, multimodal, multi-action, and comprehensive fitness action quality dataset. FLEX leverages a high-precision markerless motion capture system to record 20 weight-loaded fitness actions performed by 38 subjects across three ability levels, capturing RGB video from five viewpoints, 3D pose data, surface EMG signals, and physiological parameters. Moreover, FLEX is the first AQA dataset to integrate sEMG and a knowledge graph: by defining penalty-function-based annotation rules, we establish correspondences among weight-loaded actions, action knots, error types, and feedback. We conduct baseline experiments and novel applications on FLEX to demonstrate the effectiveness and potential of its multimodal data. We anticipate that FLEX will serve as a high-quality, extensible resource for researchers in the field.
Acknowledgements
This work is supported by the Suzhou Basic Scientific Research Project, Grant number SSD2024013. Fast Move Technology is especially acknowledged for providing sEMG devices, which greatly contributed to this research.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] Please refer to section 1.
    2. (b)
      Did you describe the limitations of your work? [Yes] Please refer to section 5.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] Please refer to the Appendix.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Please refer to the Appendix.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Please refer to the Appendix.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] We did not report error bars due to resource constraints, although we followed the standard evaluation setup commonly used in AQA studies.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] Please refer to the Appendix.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] Please refer to section 4.
    2. (b)
      Did you mention the license of the assets? [Yes] Please refer to the Appendix.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] Please refer to the Appendix. And we got the permission from chair that we put the full version of the dataset on self self-hosted cloud drive.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] Please refer to section 3.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] Please refer to section 3.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] Please refer to the Appendix.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [Yes] Please refer to section 3
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] Please refer to the Appendix.
7 FLEX Dataset
7.1 Action/Exercise Selection
We began by analyzing action types by conducting a comprehensive survey of publicly available fitness AQA datasets[52]. Our survey revealed that the Waseda-Squat[8] and Egoexo-Fitness[7] comprise a total of 12 self-loaded fitness actions, whereas the Fitness-AQA[6] includes 3 weight-loaded fitness actions. Self-loaded fitness actions do not rely on equipment, making them simpler and less prone to injury—thus more suitable for limited spaces or beginners. Conversely, weight-loaded fitness actions typically involve greater difficulty, with clearly defined target muscle groups and structured progression paths (e.g., systematically increasing dumbbell or barbell weights). Given the relatively narrow range of weight-loaded fitness actions in existing datasets, the FLEX dataset prioritizes these actions. However, due to time and resource constraints, collecting every possible weight-loaded fitness action was not feasible. Therefore, based on three dimensions—action prevalence, target muscle groups, and required equipment - we ultimately selected 20 common weight-loaded actions (including 10 barbell-based and 10 dumbbell-based actions). The chosen fitness actions not only closely mirror practical training regimens but also comprehensively cover both upper and lower body muscle groups, offering a more complete dataset compared to Egoexo-Fitness[7] (see in Figure 9 and Table 2). Additionally, researchers at Beijing Sport University validated the placement locations of the sEMG sensors on the target muscles for each action.
Figure 9: The overview of the FLEX actions.
Table 2: Comparison between the FLEX dataset with Egoexo-Fitness regarding detailed action types and target muscles. Both Egoexo-Fitness and FLEX are AQA datasets within the fitness domain. While Egoexo-Fitness primarily focuses on actions utilizing an individual’s body weight, FLEX specifically emphasizes fitness actions involving external equipment-based loads. In comparison, FLEX covers actions characterized by greater complexity and targets a more comprehensive range of muscle groups.
Dataset Action Type Target Muscle
Egoexo-Fitness[7] 1. Kneeling Push-ups 2. Push-ups 1. Pectoralis major 2. Anterior deltoid
3. Kneeling Torso Twist 4. Knee Raise + Abs Contract 3. Triceps brachii 4. External obliques
5. Shoulder Bridge 6. Sit-ups 5. Internal obliques 6. Rectus abdominis
7. Leg Reverse Lunge 8. Leg Lunge with Knee Lift 7. Iliopsoas 8. Gluteus maximus
9. Sumo Squat 10. Jumping Jacks 9. Hamstrings 10. Erector spinae
11. High Knee 12. Clap Jacks 11. Quadriceps 12. Adductors
13. Hip abductors
FLEX 1. Standing Barbell Overhead Press 2. Barbell Bicep Curl 1. Pectoralis major 2. Anterior deltoid
3. Barbell Upright Row 4. Dumbbell Front Raise 3. Middle deltoid 4. Posterior deltoid
5. Dumbbell Bicep Curl 6. Dumbbell Lateral Raise 5. Triceps brachii 6. Biceps brachii
7. Bent-Over Dumbbell Reverse Fly 8. Flat Barbell Bench Press 7. Brachialis 8. Supraspinatus
9. Incline Barbell Bench Press 10. Decline Barbell Bench Press 9. Trapezius 10. Rhomboids
11. Flat Dumbbell Bench Press 12. Incline Dumbbell Bench Press 11. Gluteus maximus 12. Quadriceps
13. Dumbbell Fly 14. Seated Dumbbell Shoulder Press 13. Hamstrings 14. Erector spinae
15. Deadlift 16. Squat 15. Latissimus dorsi
17. Barbell Bent-Over Row 18. T-Bar Row
19. Dumbbell Bent-Over Row 20. Dumbbell Bent-Over Triceps Extension
7.2 Subject Recruitment
Humans, as the core component in action performance, directly influence action quality. To collect more comprehensive data, the FLEX dataset required subjects across various capability levels compared with datasets that only contained professional-level subjects. So, we extensively recruited subjects within our institution and local commercial gyms, and ultimately selected 38 male subjects, comprising 10 professional coaches, 8 amateurs, and 20 novices. We entered into a data-collection agreement with each subject, securing their authorization to record video and physiological data and to make these publicly available for academic research; the agreement also explicitly delineates the applicable financial compensation terms. As noted above, to ensure data quality, only the professional group was required to complete all 20 action tests in one continuous session; the remaining 28 subjects were divided by target muscle groups and completed their tests incrementally over several days. Given the differences in testing intensity, we implemented a differentiated compensation scheme: professional subjects received ¥350 for completing all tests within approximately two hours, while the other subjects completed their sessions across three days, about 30 minutes per day, and were compensated ¥160 per day. (The minimum pay in Suzhou is ¥24/h.)
7.3 Collection Instruction
As noted above, prior to data collection, the experimenter outfitted each subject with a custom vest and had surface EMG sensors affixed to the target muscle regions. Subjects then entered the capture area individually, and upon the experimenter’s start signal, each subject performed the action ten times based on their own experience. Notably, the experimenter informed subjects of the specific action to be recorded before each trial, and the equipment load was set to 80% of the subject’s maximum to ensure that the action was executed by their own capability.
7.4 Annotation Rules
Leveraging established motion-analysis standards while addressing the practical demands of large-scale training data, we formulated a principled annotation protocol governed by multiple criteria. Each captured exercise instance is decomposed into three temporal phases, defined by dominant limb kinematics: Preparation, Concentric, and Eccentric. Within every phase, the primary limb movement is further factorised into an ordered sequence of action key-steps (AKs), each paired with a concise semantic description. Concatenating these descriptions produces a textual narrative of the complete exercise.
We also catalogue fine-grained error types (ETs) that characterise local joint deviations and specify prescriptive feedback for every major limb action. These entities—actions, AKs, ETs, feedback messages, and their relations—are encoded in a Fitness-Action Knowledge Graph (see Figure 10). To ensure rapid yet consistent labelling, we introduce a penalty-based scoring scheme in which distinct ETs carry heterogeneous weights; annotators simply select the observed ETs in a clip, and the overall score is obtained by summing their penalties. LABEL:tablex_all_rules and LABEL:tablex_et_weight enumerate the complete annotation rule set and the associated ET weights, respectively.
The criteria for assigning weights to error types are as follows:
  • Errors that directly compromise joint safety (e.g., spine, knee, hip, shoulder) receive the highest weight (9–10).
  • Errors that undermine core stability, disrupt movement trajectory, or alter force generation (e.g., relying on momentum) receive high weight (7–9).
  • Errors in foundational posture (e.g., initial stance, foot spacing) and insufficient movement amplitude receive moderate weight (6–7).
  • Auxiliary errors related to fine control receive a lower weight (5).
  • Breathing errors (incorrect inhalation or exhalation) receive the minimal weight (2).
Figure 10: The overview of the FLEX knowledge graph. (a) Visualization of frequently used annotation words. (b) FLEX-KG: the structure of the knowledge graph. (c1) Mapping between actions and action knots. (c2) Mapping between action knots and error types.
7.5 Annotator Recruitment
Due to the vast volume of the FLEX dataset and stringent annotation quality requirements, we recruited 16 professional practitioners from the fitness domain, including fitness trainers and graduate students in sports science, to participate in the annotation work. To ensure consistency and scientific rigor throughout the annotation process, we engaged in in-depth collaboration with all annotators before project commencement, clarifying the annotation requirements, personnel qualifications, and the underlying logic of the guidelines. To validate the feasibility and authority of these rules, we initially conducted a small-scale pilot annotation, during which we reviewed the annotators’ professional credentials and, based on the pilot results, made necessary revisions to the guidelines and refined the selection of annotators. The payment of a single annotator is higher than the minimum pay in Suzhou.
7.6 Score Calculation
During the annotation-rule formulation phase, we assigned different penalty weights to error types according to their impact on action quality. We then summed the weights of all error types present in a given action to obtain that action’s total weight. For each sample, based on the error types identified by the annotators, we calculated the ratio of the sample’s cumulative error weight to the action’s total weight, and subsequently computed the action score using the following formula (see in Figure 11):
$Score=[1-\frac{\sum_{i=1}^{N}W_{i}}{\sum_{j=1}^{M}W_{j}}]\times 100$ (1)
where $N$ represents the number of errors the sample contains, $W_{i}$ represents the weight of errors that the sample contains, $M$ represents the number of errors that the action type contains, and $W_{j}$ represents the weight of errors that the action type contains.
Figure 11: Visualization of exemplary errors and scoring during one of the exercises—barbell overhead press. Several key errors were observed that could compromise form and effectiveness. First, in the preparation process, the stance was too narrow and the grip was open rather than closed. When pressing overhead, trunk swaying and excessive elbow hyperextension were noted. The barbell was lowered too quickly, while excessive forward spinal curvature was present. Based on these errors, the final score for this action was computed to be 77.02.
7.7 Data Statistics
FLEX dataset comprises 7600 experimental trials in which 38 subjects across three ability levels performed 20 distinct weight-loaded fitness actions. After data cleaning and annotation verification, 7,512 action samples were retained. We then conducted statistical analyses of per-action sample counts, sample durations, and action quality scores, shown in Figure 12. Each action’s sample number is uniformly distributed between 370 and 380. The overall average duration is 233.76 frames, with per-action durations spanning from 100 to 1,000 frames. The dataset’s mean quality score is 68.15, exhibiting an approximately normal distribution, with per-action quality scores ranging from 20 to 100. We also report the top 20 most frequent error types. Additionally, we provide the weight loads each subject uses for each action Figure 13.
Figure 12: (a) Sample number of 20 actions. (b) Average duration and score of 20 actions. (c) Overall score distribution of the dataset and the top 20 most frequent error types.
Figure 13: Weight of subjects loaded in fitness actions. FLEX comprises 20 weight-loaded fitness actions evenly divided between barbells and dumbbells. For barbells, the intrinsic weight of the bar (20 kg) is included in the calculation, whereas for dumbbells, only the single-sided weight is considered. In the figure, the X-axis denotes different subjects, and the Y-axis indicates the various actions. The color intensity reflects differences in weight magnitude, with the darker one corresponding to heavier weights.
7.8 Data Protection
The FLEX dataset will be made available for download to all researchers. However, to protect subject privacy, all facial regions in the figures presented in this paper have been blurred, and researchers must sign a License Agreement before receiving the download link, thereby ensuring that the data are used exclusively for non-commercial academic research and that personal privacy is rigorously safeguarded.
8 Experiments
8.1 Action Quality Assessment
CoRe[48] consists of an I3D[53] pre-trained on the Kinetics dataset and a Group-Aware Regression Tree (GART). The processing pipeline is as follows: for each video pair (exemplar and target), 1×1024-dimensional features are extracted from both videos using I3D. The target’s ground-truth quality score (1×1) is then concatenated with the two feature vectors to form a 1×2049-dimensional pair feature, which is passed to GART. The GART performs a coarse-to-fine procedure: it first classifies the pair feature into groups and then regresses the score difference between the exemplar and the target. The predicted quality score for the target is obtained by adding this difference to the exemplar’s known score.
Setup.
To quantify the benefit of multimodal inputs for action-quality assessment, we retain the original GART architecture, loss function, and training strategy, and only substitute the pair feature input to create three CoRe variants:
Implementation.
Following CoRe’s protocol, we first segment each full action video into a single-action clip based on keyframes, then sample 103 frames per clip. Because skeletal and sEMG signals are synchronized with the video, we segment them using the same keyframe indices: each skeleton sample is represented by a sequence of 103 frames, while sEMG retains its native sequence length. We set the learning rate of the pre-trained feature extractors (I3D and ST-GCN) to $1\times 10^{-4}$ and the GART learning rate to $1\times 10^{-3}$, optimizing with Adam and no weight decay. Unlike the original CoRe, which uses 10 voters on MTL-AQA and AQA-7, we employ 3 on FLEX. Experiments were conducted on multiple NVIDIA RTX 4090 and A40 GPUs; each model was trained for 200 epochs per action, consuming roughly 1,600 GPU-hours.
Metrics.
In our AQA experiments, we use Spearman’s rank correlation ($\rho$) and Relative $L2$-Distance ($R\mbox{-}l2$) as evaluation metrics. $\rho$, introduced by Pirsiavash et al.[1], is the most widely adopted performance indicator in the AQA field; it quantifies the strength and direction of the monotonic relationship between ground-truth and predicted rankings but does not capture the absolute differences in their scores. The $\rho$ formula is as follows:
$\rho=1-\frac{6\sum_{i=1}^{n}(R_{i}-\widehat{R_{i}})^{2}}{n(n^{2}-1)}$ (2)
where $R_{i}$ and $\widehat{R}_{i}$ represent the ground truth and predicted rankings of the $i$-th sample, respectively. $n$ is the total number of samples. The $\rho$ range is $[-1,1]$, with values closer to 1 indicating better performance. SRC is currently the most widely used performance metric in AQA. However, SRC can only measure the strength and direction of the monotonic relationship between ground truth and predicted rankings without measuring the differences between ground truth and predicted scores.
To address this limitation, Yu et al.[48]proposed $R\mbox{-}l2$, an $L2$-based metric that, unlike $\rho$, emphasizes the numerical discrepancy between ground-truth and predicted scores. Furthermore, by normalizing for the score ranges of different action categories, $R\mbox{-}l2$ facilitates cross-category training in a way that traditional $L2$ distance cannot. The $R-l2$ formula is as follows:
$R-l2=\frac{1}{n}\sum_{i=1}^{n}(\frac{|s_{i}-\widehat{s_{i}}|}{s_{max}-s_{min}})^{2}$ (3)
where $s_{i}$ and $\widehat{s_{i}}$ represent the ground truth and predicted scores of the $i$-th sample. $s_{max}$ and $s_{min}$ represent the maximum and minimum scores of this action category. $n$ is the total number of samples. The $R-l2$ range is $[0,1]$, closer to 0 indicating better performance.
8.2 Action Quality Understanding
Implementation.
For Qwen2.5-VL-SFT, supervised fine-tuning was conducted using llama-factory[54]. We used bfloat16 precision, LoRA with rank = 8, alpha = 16, and a dropout rate of 0.05. The optimizer was AdamW with an initial learning rate of $5\times 10^{-5}$, cosine-decay scheduling, and gradient clipping at 1.0. The batch size was 2 with gradient accumulation over 4 steps, and training spanned 2 epochs for 12 hours on a single NVIDIA H20.
Figure 14: The construction of FLEX-VideoQA dataset. We designed a dialogue template following the pipeline “action recognition → action standards → action evaluation → action scoring,” with all questions and reference answers automatically generated from our annotation rules and results. In particular, action-evaluation answers were pre-generated by DeepseekV3 by combining video samples, action knots, error types, and feedback suggestions to ensure consistency.
8.3 Video2EMG
Implementation.
Image data were prepared according to CoRe’s specifications; the key difference is that we applied a sliding-window average to the sEMG signals to be temporally aligned per frame with the video. We used PyTorch’s pre-trained ResNet-50 as our feature extractor, generating a 2,048-dimensional feature vector for each frame. These vectors were then passed through a lightweight convolutional layer followed by a single-layer, unidirectional LSTM to predict the corresponding sEMG signal. The model was trained with mean squared error (MSE) loss, using the Adam optimizer (initial learning rate $1\times 10^{-3}$), a batch size of 32, and 150 epochs for 6 hours on a single NVIDIA RTX4090. Meanwhile, we trained and tested the model’s performance on 18 different muscles according to the sEMG collected (see in Table 3).
Table 3: Muscles Recorded by the sEMG Device in the FLEX Dataset.
Action Muscle
Standing Barbell Overhead Press Left anterior deltoid Right anterior deltoid Left medial deltoid Right medial deltoid
Barbell Biceps Curl Left biceps brachii Right biceps brachii Left triceps brachii Right triceps brachii
Barbell Upright Row Left medial deltoid Right medial deltoid Left anterior deltoid Right anterior deltoid
Dumbbell Front Raise Left anterior deltoid Right anterior deltoid Left medial deltoid Right medial deltoid
Dumbbell Biceps Curl Left biceps brachii Right biceps brachii Left triceps brachii Right triceps brachii
Dumbbell Lateral Raise Left medial deltoid Right medial deltoid Left anterior deltoid Right anterior deltoid
Bent-Over Dumbbell Reverse Fly Left posterior deltoid Right posterior deltoid
Flat Barbell Bench Press Left pectoralis major Right pectoralis major Left pectoralis major Right pectoralis major
Incline Barbell Bench Press Left pectoralis major Right pectoralis major Left pectoralis major Right pectoralis major
Decline Barbell Bench Press Left pectoralis major Right pectoralis major Left pectoralis major Right pectoralis major
Flat Dumbbell Bench Press Left pectoralis major Right pectoralis major Left pectoralis major Right pectoralis major
Incline Dumbbell Bench Press Left pectoralis major Right pectoralis major Left pectoralis major Right pectoralis major
Dumbbell Fly Left pectoralis major Right pectoralis major Left pectoralis major Right pectoralis major
Seated Dumbbell Shoulder Press Left anterior deltoid Right anterior deltoid Left medial deltoid Right medial deltoid
Deadlift Left biceps femoris Right biceps femoris Left vastus lateralis Right vastus lateralis
Squat Left vastus lateralis Right vastus lateralis Left biceps femoris Right biceps femoris
Barbell Bent-Over Row Left latissimus dorsi Right latissimus dorsi
T-Bar Row Left latissimus dorsi Right latissimus dorsi
Dumbbell Bent-Over Row Left latissimus dorsi Right latissimus dorsi
Dumbbell Bent-Over Triceps Extension Left triceps brachii Right triceps brachii Left biceps brachii Right biceps brachii
9 Data Availability
The FLEX dataset’s raw data volume approaches 3 TB, exceeding the storage capacity of the official platforms NeurIPS recommended; with approval from the D&B Chair, we have deployed it on self-hosted cloud storage. Dataset and code are released under the CC BY-NC 4.0 license.
10 Potential Negative Impact
The capability of measuring the quality of actions and movements may be used in ways that users/humans may not approve/agree. For example, action quality assessment may be misused by agencies such as health insurance providers to forecast future injuries and diseases reliably. Such health insurance providers may not then unjustly cover future injury-prone clientele. Awareness is the first step towards the solution. We believe our work helps create this awareness. Moreover, detecting bias of this nature in the decision-making process can help provide more equitable service and consideration to society.
11 Croissant Checker
The dataset’s Croissant file has been validated using the official NeurIPS tool, and the validation results are shown in the Figure 15 below.
Figure 15: The result of croissant checker.
Table 4: Full list of the FLEX annotation rules.
Standing Barbell Overhead Press Action Keystep Error Type
Preparation AK01: Stand with feet shoulder-width apart and toes slightly turned outward. Keep the chest lifted and the core engaged. Maintain a neutral spine position with the scapulae depressed and retracted, ensuring trunk stability. ET001: Feet positioned too narrowly or too widely.
ET002: Excessive arching or collapsing of the lower back.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
AK02: Grip the barbell with a pronated (overhand) closed grip, with hands slightly wider than shoulder-width apart. Keep the elbows slightly flexed and maintain a neutral wrist position. ET006: Open grip.
ET007: Grip width is too narrow or too wide.
ET008: Excessive elbow flexion or hyperextension.
ET009: Wrist excessively extended or tilted to the side.
AK03: Lift the barbell to rest above the clavicles. Forearms remain perpendicular to the floor, palms facing up, elbows lower than the shoulders for stable support. ET010: The initial barbell position is set too high or too low.
ET011: Forearms are not perpendicular to the ground.
ET012: Palms facing inward or downward.
ET013: Excessive elbow adduction or abduction.
Concentric AK04: Exhale. Primarily activate the anterior deltoid, assisted by the triceps, with elbows pointing forward. Press the barbell overhead until the arms are straight. Maintain a neutral spine and braced core throughout. ET014: Inhalation.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET005: Trunk swaying forward and backward, indicating instability.
ET008: Excessive elbow flexion or hyperextension.
ET013: Excessive elbow adduction or abduction.
ET015: Using a knee drive or leaning the torso backward to assist the movement.
ET016: Excessive forward head movement, obstructing the barbell’s vertical path.
Eccentric AK05: Inhale. Control the barbell along the same path, lowering it smoothly to the clavicles. Keep the scapulae stable, the deltoids resisting gravity, and forearms vertical. Engage the core and maintain a neutral lumbar spine. ET017: Exhalation.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET011: Forearms are not perpendicular to the ground.
ET013: Excessive elbow adduction or abduction.
ET018: Lowering too quickly.
ET019: Ending position of the barbell is too high or too low.
Barbell Biceps Curl Action Keystep Error Type
Preparation AK01: Stand with feet shoulder-width apart and toes slightly turned outward. Keep the chest lifted and the core engaged. Maintain a neutral spine position with the scapulae depressed and retracted, ensuring trunk stability. ET001: Feet positioned too narrowly or too widely.
ET002: Excessive arching or collapsing of the lower back.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
AK06: Grip the barbell with an underhand, closed grip, slightly wider than shoulder-width. Keep the wrists in a neutral position. ET006: Open grip.
ET007: Grip width is too narrow or too wide.
ET009: Wrist excessively extended or tilted to the side.
AK07: Arms hang naturally with elbows slightly bent. Elbows remain tucked in, upper arms fixed. The barbell starts at the front of the thighs. ET008: Excessive elbow flexion or hyperextension.
ET020: Elbows move forward or backward, upper arms are not stabilized.
ET021: Excessive elbow abduction, moving away from the torso.
Concentric AK08: Exhale. Actively contract the biceps to curl the forearms along a semicircular path up to the front of the shoulders. Apply symmetrical force with both hands to maintain balance. Keep the upper arms fixed and spine neutral throughout. ET014: Inhalation.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET005: Trunk swaying forward and backward, indicating instability.
ET009: Wrist excessively extended or tilted to the side.
ET015: Using a knee drive or leaning the torso backward to assist the movement.
ET019: Ending position of the barbell is too high or too low.
ET022: Using forward or upward shoulder movement to assist the action.
ET023: Elbows move forward or backward, causing the barbell path to deviate from an arc.
Eccentric AK09: Inhale. The biceps control the barbell on the slow descent back to the starting position at the front of the thighs. Keep the elbows slightly bent (not locked) and the upper arms tight against the torso. ET017: Exhalation.
ET008: Excessive elbow flexion or hyperextension.
ET018: Lowering too quickly.
ET020: Elbows move forward or backward, upper arms are not stabilized.
Barbell Upright Row Action Keystep Error Type
Preparation AK01: Stand with feet shoulder-width apart and toes slightly turned outward. Keep the chest lifted and the core engaged. Maintain a neutral spine position with the scapulae depressed and retracted, ensuring trunk stability. ET001: Feet positioned too narrowly or too widely.
ET002: Excessive arching or collapsing of the lower back.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
AK02: Grip the barbell with a pronated (overhand) closed grip, with hands slightly wider than shoulder-width apart. Keep the elbows slightly flexed and maintain a neutral wrist position. ET006: Open grip.
ET007: Grip width is too narrow or too wide.
ET008: Excessive elbow flexion or hyperextension.
ET009: Wrist excessively extended or tilted to the side.
AK10: Position the barbell in front of the thighs with a firm grip, ensuring the web of the hands is securely locked around the bar. ET010: The initial barbell position is set too high or too low.
Concentric AK11: Exhale. Maintain a stable torso. Focus on contracting the middle deltoid. Abduct the elbows out to the sides and lift the barbell vertically along the sternum until it reaches chest height. ET014: Inhalation.
ET005: Trunk swaying forward and backward, indicating instability.
ET015: Using a knee drive or leaning the torso backward to assist the movement.
ET019: Ending position of the barbell is too high or too low.
ET024: Elbows rising much lower than the shoulder joints.
ET025: The barbell does not travel vertically relative to the ground.
Eccentric AK12: Inhale. Use the deltoids’ eccentric strength to control the barbell’s descent along the same path back to the start. Keep the lumbar spine neutral and core braced, with the elbows slightly bent at the bottom. ET017: Exhalation.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET008: Excessive elbow flexion or hyperextension.
ET018: Lowering too quickly.
Dumbbell Front Raise Action Keystep Error Type
Preparation AK01: Stand with feet shoulder-width apart and toes slightly turned outward. Keep the chest lifted and the core engaged. Maintain a neutral spine position with the scapulae depressed and retracted, ensuring trunk stability. ET001: Feet positioned too narrowly or too widely.
ET002: Excessive arching or collapsing of the lower back.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
AK13: Arms fully extended. Hold the dumbbells in an overhand, closed grip at the front of the thighs, palms facing the legs, elbows slightly bent. ET006: Open grip.
ET008: Excessive elbow flexion or hyperextension.
ET026: Palms facing forward or backward.
Concentric AK14: Exhale. Focus on the anterior deltoid as the primary mover. Using the shoulder joint as the pivot, smoothly raise the dumbbells forward until the elbows are level with the shoulders. Keep the core stable, shoulders depressed, and elbows slightly bent. ET014: Inhalation.
ET005: Trunk swaying forward and backward, indicating instability.
ET008: Excessive elbow flexion or hyperextension.
ET015: Using a knee drive or leaning the torso backward to assist the movement.
ET022: Using forward or upward shoulder movement to assist the action.
ET027: Changing the elbow angle, causing the dumbbell path to deviate from an arc.
ET028: Ending position of the dumbbell is too high or too low.
Eccentric AK15: Inhale. Maintain tension in the deltoids to resist gravity. Control the dumbbells as you lower them along the same path back to the starting position. Keep the core stable and elbows slightly bent.
ET017: Exhalation.
ET008: Excessive elbow flexion or hyperextension.
ET018: Lowering too quickly.
Dumbbell Biceps Curl Action Keystep Error Type
Preparation AK01: Stand with feet shoulder-width apart and toes slightly turned outward. Keep the chest lifted and the core engaged. Maintain a neutral spine position with the scapulae depressed and retracted, ensuring trunk stability. ET001: Feet positioned too narrowly or too widely.
ET002: Excessive arching or collapsing of the lower back.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
AK16: Grip the dumbbells with a neutral, closed grip. ET006: Open grip.
AK07: Arms hang naturally with elbows slightly bent. Elbows remain tucked in, upper arms fixed. The barbell starts at the front of the thighs. ET008: Excessive elbow flexion or hyperextension.
ET020: Elbows move forward or backward, upper arms are not stabilized.
ET021: Excessive elbow abduction, moving away from the torso.
Concentric AK17: Exhale. Actively contract the biceps, curling the forearms in a semicircle while externally rotating the forearms, until the dumbbells reach shoulder level. Keep the upper arms fixed and the spine neutral. ET014: Inhalation.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET005: Trunk swaying forward and backward, indicating instability.
ET015: Using a knee drive or leaning the torso backward to assist the movement.
ET022: Using forward or upward shoulder movement to assist the action.
ET028: Ending position of the dumbbell is too high or too low.
ET029: Elbows move forward or backward, causing the dumbbell path to deviate from an arc.
ET030: Excessive wrist supination over-rotating the forearms.
Eccentric AK18: Inhale. The biceps control the dumbbells on the slow descent back to the starting position in front of the thighs. Keep elbows slightly bent and upper arms close to the torso. ET017: Exhalation.
ET008: Excessive elbow flexion or hyperextension.
ET018: Lowering too quickly.
ET020: Elbows move forward or backward, upper arms are not stabilized.
Dumbbell Lateral Raise Action Keystep Error Type
Preparation AK01: Stand with feet shoulder-width apart and toes slightly turned outward. Keep the chest lifted and the core engaged. Maintain a neutral spine position with the scapulae depressed and retracted, ensuring trunk stability. ET001: Feet positioned too narrowly or too widely.
ET002: Excessive arching or collapsing of the lower back.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
AK19: Hold the dumbbells in a neutral, closed grip, positioning them in front of the body with arms fully extended. ET006: Open grip.
ET008: Excessive elbow flexion or hyperextension.
ET031: Initial position of the dumbbell is too far forward or backward.
Concentric AK20: Exhale. Focus on the anterior and middle deltoid. Using the shoulder joint as the pivot, lift the dumbbells laterally to shoulder height. Keep elbows slightly bent, core stable, and shoulders depressed throughout. ET014: Inhalation.
ET005: Trunk swaying forward and backward, indicating instability.
ET008: Excessive elbow flexion or hyperextension.
ET015: Using a knee drive or leaning the torso backward to assist the movement.
ET022: Using forward or upward shoulder movement to assist the action.
ET027: Changing the elbow angle, causing the dumbbell path to deviate from an arc.
ET028: Ending position of the dumbbell is too high or too low.
Eccentric AK15: Inhale. Maintain tension in the deltoids to resist gravity. Control the dumbbells as you lower them along the same path back to the starting position. Keep the core stable and elbows slightly bent.
ET017: Exhalation.
ET008: Excessive elbow flexion or hyperextension.
ET018: Lowering too quickly.
Bent-Over Dumbbell Reverse Fly Action Keystep Error Type
Preparation AK21: Stand with feet shoulder-width apart, toes slightly turned outward, and knees slightly flexed. Keep the back straight and bend forward to approximately a 40° angle with the ground. ET001: Feet positioned too narrowly or too widely.
ET002: Excessive arching or collapsing of the lower back.
ET032: Excessive knee flexion or hyperextension.
ET033: Torso angle too shallow or too steep.
AK22: Let the arms hang naturally, holding the dumbbells with a neutral, closed grip just below the knee level. Keep the elbows slightly flexed.
ET006: Open grip.
ET008: Excessive elbow flexion or hyperextension.
Concentric AK23: Exhale. Focus on the posterior deltoid. Using the shoulder joint as the pivot, lift the dumbbells laterally until they reach shoulder height. Keep elbows slightly bent, core stable, and shoulders depressed. ET014: Inhalation.
ET008: Excessive elbow flexion or hyperextension.
ET027: Changing the elbow angle, causing the dumbbell path to deviate from an arc.
ET034: During abduction, the scapula elevates and the neck protrudes forward.
ET035: Arms extending diagonally backward, lifting beyond the shoulder joint’s horizontal level.
Eccentric AK15: Inhale. Maintain tension in the deltoids to resist gravity. Control the dumbbells as you lower them along the same path back to the starting position. Keep the core stable and elbows slightly bent.
ET017: Exhalation.
ET008: Excessive elbow flexion or hyperextension.
ET018: Lowering too quickly.
Flat Barbell Bench Press Action Keystep Error Type
Preparation AK24: Lie supine on a flat bench, feet flat on the floor. Keep the waist and abdomen naturally braced, shoulders depressed, and ensure the head, upper back, and hips remain on the bench. Lift the chest and engage the abdomen. ET036: Feet not firmly planted on the ground.
ET002: Excessive arching or collapsing of the lower back.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
ET037: Head, upper back, or glutes lifting off the bench.
AK25: Fully extend the arms, gripping the barbell with an overhand, closed grip, slightly wider than shoulder-width. ET006: Open grip.
ET007: Grip width is too narrow or too wide.
ET009: Wrist excessively extended or tilted to the side.
Eccentric AK26: Inhale. Maintain torso stability, allow the shoulders to naturally depress, and bend the elbows downward so the barbell lowers slowly and smoothly to the chest or just above it. ET017: Exhalation.
ET018: Lowering too quickly.
ET019: Ending position of the barbell is too high or too low.
ET038: The angle between the upper arm and torso is too large or too small.
ET039: Insufficient barbell descent, resulting a shortened range of motion.
Concentric AK27: Exhale. Focus on pectoralis major contraction. Using the shoulder joint as the pivot, apply symmetrical force with both hands to press the barbell upward. Keep the elbows slightly bent at the top. ET014: Inhalation.
ET008: Excessive elbow flexion or hyperextension.
ET040: Hips rise or legs drive off the floor to assist during the press.
ET041: Raising the head or pushing the neck forward to assist during the press.
ET042: Shrugging the shoulders or shifting forward during the press.
ET043: The barbell sways left or right, or one arm dominates the force output.
ET044: Insufficient barbell elevation, resulting a shortened range of motion.
Incline Barbell Bench Press Action Keystep Error Type
Preparation AK28: Lie supine on an incline bench at approximately 30° to the floor, feet flat. Keep the waist and abdomen braced, shoulders depressed, and the head, upper back, and hips on the bench. Chest lifted and abdomen engaged. ET002: Excessive arching or collapsing of the lower back.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
ET036: Feet not firmly planted on the ground.
ET037: Head, upper back, or glutes lifting off the bench.
AK25: Fully extend the arms, gripping the barbell with an overhand, closed grip, slightly wider than shoulder-width. ET006: Open grip.
ET007: Grip width is too narrow or too wide.
ET009: Wrist excessively extended or tilted to the side.
Eccentric AK26: Inhale. Maintain torso stability, allow the shoulders to naturally depress, and bend the elbows downward so the barbell lowers slowly and smoothly to the chest or just above it. ET017: Exhalation.
ET018: Lowering too quickly.
ET019: Ending position of the barbell is too high or too low.
ET038: The angle between the upper arm and torso is too large or too small.
ET039: Insufficient barbell descent, resulting a shortened range of motion.
Concentric AK27: Exhale. Focus on pectoralis major contraction. Using the shoulder joint as the pivot, apply symmetrical force with both hands to press the barbell upward. Keep the elbows slightly bent at the top. ET014: Inhalation.
ET008: Excessive elbow flexion or hyperextension.
ET040: Hips rise or legs drive off the floor to assist during the press.
ET041: Raising the head or pushing the neck forward to assist during the press.
ET042: Shrugging the shoulders or shifting forward during the press.
ET043: The barbell sways left or right, or one arm dominates the force output.
ET044: Insufficient barbell elevation, resulting a shortened range of motion.
Decline Barbell Bench Press Action Keystep Error Type
Preparation AK29: Lie supine on a decline bench at approximately 15° to the floor, feet flat. Keep the waist and abdomen braced, shoulders depressed, and the head, upper back, and hips on the bench. Chest lifted and abdomen engaged. ET002: Excessive arching or collapsing of the lower back.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
ET036: Feet not firmly planted on the ground.
ET037: Head, upper back, or glutes lifting off the bench.
AK25: Fully extend the arms, gripping the barbell with an overhand, closed grip, slightly wider than shoulder-width. ET006: Open grip.
ET007: Grip width is too narrow or too wide.
ET009: Wrist excessively extended or tilted to the side.
Eccentric AK26: Inhale. Maintain torso stability, allow the shoulders to naturally depress, and bend the elbows downward so the barbell lowers slowly and smoothly to the chest or just above it. ET017: Exhalation.
ET018: Lowering too quickly.
ET019: Ending position of the barbell is too high or too low.
ET038: The angle between the upper arm and torso is too large or too small.
ET039: Insufficient barbell descent, resulting a shortened range of motion.
Concentric AK27: Exhale. Focus on pectoralis major contraction. Using the shoulder joint as the pivot, apply symmetrical force with both hands to press the barbell upward. Keep the elbows slightly bent at the top. ET014: Inhalation.
ET008: Excessive elbow flexion or hyperextension.
ET040: Hips rise or legs drive off the floor to assist during the press.
ET041: Raising the head or pushing the neck forward to assist during the press.
ET042: Shrugging the shoulders or shifting forward during the press.
ET043: The barbell sways left or right, or one arm dominates the force output.
ET044: Insufficient barbell elevation, resulting a shortened range of motion.
Flat Dumbbell Bench Press Action Keystep Error Type
Preparation AK24: Lie supine on a flat bench, feet flat on the floor. Keep the waist and abdomen naturally braced, shoulders depressed, and ensure the head, upper back, and hips remain on the bench. Lift the chest and engage the abdomen. ET036: Feet not firmly planted on the ground.
ET002: Excessive arching or collapsing of the lower back.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
ET037: Head, upper back, or glutes lifting off the bench.
AK30: Hold the dumbbells with a neutral, closed grip. Bend the elbows, keeping the palms facing forward and forearms perpendicular to the ground, supporting the dumbbells on either side of the pectoralis major. ET006: Open grip.
ET009: Wrist excessively extended or tilted to the side.
ET011: Forearms are not perpendicular to the ground.
ET045: Initial position of the dumbbell is too high or too low.
Concentric AK31: Exhale. Maintain torso stability. Focus on contracting the pectoralis major to press both dumbbells inward and upward until the arms are extended above the chest. ET014: Inhalation.
ET008: Excessive elbow flexion or hyperextension.
ET040: Hips rise or legs drive off the floor to assist during the press.
ET041: Raising the head or pushing the neck forward to assist during the press.
ET042: Shrugging the shoulders or shifting forward during the press.
ET046: Dumbbells collide or deviate outward.
ET047: Uneven elevation of the dumbbells on both sides.
ET048: Insufficient dumbbell elevation, resulting a shortened range of motion.
Eccentric AK32: Inhale. Let the shoulders depress naturally, forearms remain vertical. Bend the elbows downward in a controlled manner until the dumbbells are at the sides of the shoulders. ET017: Exhalation.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
ET011: Forearms are not perpendicular to the ground.
ET018: Lowering too quickly.
ET038: The angle between the upper arm and torso is too large or too small.
ET049: Insufficient dumbbell descent, resulting a shortened range of motion.
Incline Dumbbell Bench Press Action Keystep Error Type
Preparation AK28: Lie supine on an incline bench at approximately 30° to the floor, feet flat. Keep the waist and abdomen braced, shoulders depressed, and the head, upper back, and hips on the bench. Chest lifted and abdomen engaged. ET002: Excessive arching or collapsing of the lower back.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
ET036: Feet not firmly planted on the ground.
ET037: Head, upper back, or glutes lifting off the bench.
AK30: Hold the dumbbells with a neutral, closed grip. Bend the elbows, keeping the palms facing forward and forearms perpendicular to the ground, supporting the dumbbells on either side of the pectoralis major. ET006: Open grip.
ET009: Wrist excessively extended or tilted to the side.
ET011: Forearms are not perpendicular to the ground.
ET045: Initial position of the dumbbell is too high or too low.
Concentric AK31: Exhale. Maintain torso stability. Focus on contracting the pectoralis major to press both dumbbells inward and upward until the arms are extended above the chest. ET014: Inhalation.
ET008: Excessive elbow flexion or hyperextension.
ET040: Hips rise or legs drive off the floor to assist during the press.
ET041: Raising the head or pushing the neck forward to assist during the press.
ET042: Shrugging the shoulders or shifting forward during the press.
ET046: Dumbbells collide or deviate outward.
ET047: Uneven elevation of the dumbbells on both sides.
ET048: Insufficient dumbbell elevation, resulting a shortened range of motion.
Eccentric AK32: Inhale. Let the shoulders depress naturally, forearms remain vertical. Bend the elbows downward in a controlled manner until the dumbbells are at the sides of the shoulders. ET017: Exhalation.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
ET011: Forearms are not perpendicular to the ground.
ET018: Lowering too quickly.
ET038: The angle between the upper arm and torso is too large or too small.
ET049: Insufficient dumbbell descent, resulting a shortened range of motion.
Dumbbell Fly Action Keystep Error Type
Preparation AK24: Lie supine on a flat bench, feet flat on the floor. Keep the waist and abdomen naturally braced, shoulders depressed, and ensure the head, upper back, and hips remain on the bench. Lift the chest and engage the abdomen. ET036: Feet not firmly planted on the ground.
ET002: Excessive arching or collapsing of the lower back.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
ET037: Head, upper back, or glutes lifting off the bench.
AK33: Hold the dumbbells with a neutral, closed grip, fully extending the arms to support the dumbbells above the chest. Keep the palms facing each other and elbows slightly flexed. ET006: Open grip.
ET008: Excessive elbow flexion or hyperextension.
ET045: Initial position of the dumbbell is too high or too low.
Eccentric AK34: Inhale. Using the pectoralis major’s strength, control the arms as they open to the sides. The lowest point is when the dumbbells align with shoulder height. ET017: Exhalation.
ET008: Excessive elbow flexion or hyperextension.
ET009: Wrist excessively extended or tilted to the side.
ET018: Lowering too quickly.
ET027: Changing the elbow angle, causing the dumbbell path to deviate from an arc.
ET049: Insufficient dumbbell descent, resulting a shortened range of motion.
ET050: Uneven descent of the dumbbells on both sides.
Concentric AK35: Exhale. Keep the shoulders depressed. Focus on the pectoralis major contraction to bring the arms smoothly back together until they are extended, with elbows slightly bent. ET014: Inhalation.
ET008: Excessive elbow flexion or hyperextension.
ET040: Hips rise or legs drive off the floor to assist during the press.
ET041: Raising the head or pushing the neck forward to assist during the press.
ET042: Shrugging the shoulders or shifting forward during the press.
ET046: Dumbbells collide or deviate outward.
ET047: Uneven elevation of the dumbbells on both sides.
ET048: Insufficient dumbbell elevation, resulting a shortened range of motion.
Seated Dumbbell Shoulder Press Action Keystep Error Type
Preparation AK36: Feet flat on the floor, upper back and hips pressed against the seat. Chest lifted, abdomen braced, maintaining a normal lumbar curve. ET002: Excessive arching or collapsing of the lower back.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
ET036: Feet not firmly planted on the ground.
ET037: Head, upper back, or glutes lifting off the bench.
AK37: Hold the dumbbells with an overhand, closed grip at shoulder level, keeping the wrists neutral and forearms perpendicular to the ground. Palms should be slightly forward, and elbows positioned in front of the shoulder joints. ET006: Open grip.
ET009: Wrist excessively extended or tilted to the side.
ET011: Forearms are not perpendicular to the ground.
ET012: Palms facing inward or downward.
ET013: Excessive elbow adduction or abduction.
Concentric AK38: Exhale. Engage the deltoids, using the shoulder joint as the pivot to press the dumbbells overhead until arms are nearly straight, elbows slightly bent at the top. ET014: Inhalation.
ET008: Excessive elbow flexion or hyperextension.
ET040: Hips rise or legs drive off the floor to assist during the press.
ET041: Raising the head or pushing the neck forward to assist during the press.
ET042: Shrugging the shoulders or shifting forward during the press.
ET046: Dumbbells collide or deviate outward.
ET047: Uneven elevation of the dumbbells on both sides.
ET048: Insufficient dumbbell elevation, resulting a shortened range of motion.
Eccentric AK39: Inhale. Bend the elbows in a controlled manner, lowering the dumbbells along the sides of the shoulders to the starting position, elbows slightly below shoulder level. ET017: Exhalation.
ET018: Lowering too quickly.
ET049: Insufficient dumbbell descent, resulting a shortened range of motion.
ET050: Uneven descent of the dumbbells on both sides.
Deadlift Action Keystep Error Type
Preparation AK01: Stand with feet shoulder-width apart and toes slightly turned outward. Keep the chest lifted and the core engaged. Maintain a neutral spine position with the scapulae depressed and retracted, ensuring trunk stability. ET001: Feet positioned too narrowly or too widely.
ET002: Excessive arching or collapsing of the lower back.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
AK40: Keep the back straight and bend the knees to squat down. Position the shins close to the barbell and grip it with an overhand, closed grip, slightly wider than shoulder-width. Ensure the knees are aligned with the direction of the toes. ET006: Open grip.
ET007: Grip width is too narrow or too wide.
ET051: Initial position of the barbell is too far forward or too far back.
ET058: The hip position is set too high or too low.
Concentric AK41: Exhale. Engage the gluteus maximus and hamstrings to extend the knees and hips, lifting the barbell off the ground. Move the hips forward until standing fully upright. ET014: Inhalation.
ET002: Excessive arching or collapsing of the lower back.
ET005: Trunk swaying forward and backward, indicating instability.
ET025: The barbell does not travel vertically relative to the ground.
ET044: Insufficient barbell elevation, resulting a shortened range of motion.
ET052: Hips rising faster than barbell, causing asynchronous force distribution between legs and hips.
ET053: Excessive forward or backward tilt of the shoulder joint.
ET062: Barbell deviating from the legs.
ET063: The knees lock out, with minimal extension in the hip and knee joints.
Eccentric AK42: Inhale. Keep the back straight, hinge at the hips by leaning forward and shifting the hips back under control to full extension range. Hips remain higher than the knees and lower than the shoulders. ET017: Exhalation.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET018: Lowering too quickly.
ET054: Knees bend prematurely, causing the barbell to circle around them.
ET062: Barbell deviating from the legs.
ET063: The knees lock out, with minimal extension in the hip and knee joints.
Squat Action Keystep Error Type
Preparation AK56: Stand with feet slightly wider than shoulder-width, toes slightly turned outward. Keep the chest lifted, engage the core, maintain a neutral spine, depress and retract the scapulae, and ensure torso stability. ET061: Feet placed too narrowly or too widely during the squat.
ET002: Excessive arching or collapsing of the lower back.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
AK43: Grip the barbell with an overhand, closed grip, wider than shoulder-width, and position it on the trapezius and deltoids. ET006: Open grip.
ET007: Grip width is too narrow or too wide.
ET010: The initial barbell position is set too high or too low.
Eccentric AK44: Inhale. Look forward, keep the back straight, flex hips and knees to descend smoothly until the angle between thighs and calves is slightly less than 90°. Ensure the knees track the direction of the toes. ET017: Exhalation.
ET002: Excessive arching or collapsing of the lower back.
ET005: Trunk swaying forward and backward, indicating instability.
ET033: Torso angle too shallow or too steep.
ET039: Insufficient barbell descent, resulting a shortened range of motion.
ET055: Excessive forward movement of the knees.
ET056: Knees caving inward.
Concentric AK45: Exhale. Keep the back straight. Primarily drive with the quadriceps to extend the hips and knees, moving the hips forward until the knees are straight or slightly bent. ET014: Inhalation.
ET002: Excessive arching or collapsing of the lower back.
ET032: Excessive knee flexion or hyperextension.
ET052: Hips rising faster than barbell, causing asynchronous force distribution between legs and hips.
Barbell Bent-Over Row Action Keystep Error Type
Preparation AK01: Stand with feet shoulder-width apart and toes slightly turned outward. Keep the chest lifted and the core engaged. Maintain a neutral spine position with the scapulae depressed and retracted, ensuring trunk stability. ET001: Feet positioned too narrowly or too widely.
ET002: Excessive arching or collapsing of the lower back.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
AK46: Keep the back straight and knees slightly flexed. Lean forward to a 30°–45° angle relative to the ground. Grip the barbell with an overhand, closed grip, slightly wider than shoulder-width, while keeping the wrists neutral. ET006: Open grip.
ET007: Grip width is too narrow or too wide.
ET009: Wrist excessively extended or tilted to the side.
ET032: Excessive knee flexion or hyperextension.
ET033: Torso angle too shallow or too steep.
AK47: Shift the hips backward and lift the barbell with straight arms to just below the knee level, keeping the elbows extended. ET008: Excessive elbow flexion or hyperextension.
ET010: The initial barbell position is set too high or too low.
ET013: Excessive elbow adduction or abduction.
Concentric AK48: Exhale. Focus on contracting the latissimus dorsi, pulling the upper arms back and elbows toward the midline of the back, drawing the barbell toward the lower abdomen. ET014: Inhalation.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
ET005: Trunk swaying forward and backward, indicating instability.
ET038: The angle between the upper arm and torso is too large or too small.
ET043: The barbell sways left or right, or one arm dominates the force output.
ET044: Insufficient barbell elevation, resulting a shortened range of motion.
ET057: Elbows flex before the scapula retracts.
Eccentric AK49: Inhale. Use the latissimus dorsi tension to control the barbell’s smooth descent until the arms are fully extended. ET017: Exhalation.
ET002: Excessive arching or collapsing of the lower back.
ET018: Lowering too quickly.
ET039: Insufficient barbell descent, resulting a shortened range of motion.
T-Bar Row Action Keystep Error Type
Preparation AK01: Stand with feet shoulder-width apart and toes slightly turned outward. Keep the chest lifted and the core engaged. Maintain a neutral spine position with the scapulae depressed and retracted, ensuring trunk stability. ET001: Feet positioned too narrowly or too widely.
ET002: Excessive arching or collapsing of the lower back.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
AK46: Keep the back straight and knees slightly flexed. Lean forward to a 30°–45° angle relative to the ground. Grip the barbell with an overhand, closed grip, slightly wider than shoulder-width, while keeping the wrists neutral. ET006: Open grip.
ET007: Grip width is too narrow or too wide.
ET009: Wrist excessively extended or tilted to the side.
ET032: Excessive knee flexion or hyperextension.
ET033: Torso angle too shallow or too steep.
AK47: Shift the hips backward and lift the barbell with straight arms to just below the knee level, keeping the elbows extended. ET008: Excessive elbow flexion or hyperextension.
ET010: The initial barbell position is set too high or too low.
ET013: Excessive elbow adduction or abduction.
Concentric AK48: Exhale. Focus on contracting the latissimus dorsi, pulling the upper arms back and elbows toward the midline of the back, drawing the barbell toward the lower abdomen. ET014: Inhalation.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
ET005: Trunk swaying forward and backward, indicating instability.
ET038: The angle between the upper arm and torso is too large or too small.
ET043: The barbell sways left or right, or one arm dominates the force output.
ET044: Insufficient barbell elevation, resulting a shortened range of motion.
ET057: Elbows flex before the scapula retracts.
Eccentric AK49: Inhale. Use the latissimus dorsi tension to control the barbell’s smooth descent until the arms are fully extended. ET017: Exhalation.
ET002: Excessive arching or collapsing of the lower back.
ET018: Lowering too quickly.
ET039: Insufficient barbell descent, resulting a shortened range of motion.
Dumbbell Bent-Over Row Action Keystep Error Type
Preparation AK01: Stand with feet shoulder-width apart and toes slightly turned outward. Keep the chest lifted and the core engaged. Maintain a neutral spine position with the scapulae depressed and retracted, ensuring trunk stability. ET001: Feet positioned too narrowly or too widely.
ET002: Excessive arching or collapsing of the lower back.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
AK50: Keep the back straight and knees slightly flexed. Lean forward to a 30°–45° angle relative to the ground. Hold the dumbbells with a neutral, closed grip while keeping the wrists neutral. ET006: Open grip.
ET007: Grip width is too narrow or too wide.
ET009: Wrist excessively extended or tilted to the side.
ET013: Excessive elbow adduction or abduction.
ET032: Excessive knee flexion or hyperextension.
ET033: Torso angle too shallow or too steep.
Concentric AK51: Exhale. Focus on the back musculature, bending the elbows to pull the dumbbells close along the sides of the body until the upper arms are parallel to the floor. ET014: Inhalation.
ET005: Trunk swaying forward and backward, indicating instability.
ET015: Using a knee drive or leaning the torso backward to assist the movement.
ET021: Excessive elbow abduction, moving away from the torso.
ET047: Uneven elevation of the dumbbells on both sides.
ET048: Insufficient dumbbell elevation, resulting a shortened range of motion.
ET057: Elbows flex before the scapula retracts.
ET058: The hip position is set too high or too low.
Eccentric AK52: Inhale. Brace the core, using the back musculature to control the dumbbells’ slow, steady descent until the arms are fully extended. ET017: Exhalation.
ET002: Excessive arching or collapsing of the lower back.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET008: Excessive elbow flexion or hyperextension.
ET018: Lowering too quickly.
ET049: Insufficient dumbbell descent, resulting a shortened range of motion.
ET050: Uneven descent of the dumbbells on both sides.
Dumbbell Bent-Over Triceps Extension Action Keystep Error Type
Preparation AK01: Stand with feet shoulder-width apart and toes slightly turned outward. Keep the chest lifted and the core engaged. Maintain a neutral spine position with the scapulae depressed and retracted, ensuring trunk stability. ET001: Feet positioned too narrowly or too widely.
ET002: Excessive arching or collapsing of the lower back.
ET003: Excessive forward flexion or excessive backward curvature of the spine.
ET004: Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders.
AK50: Keep the back straight and knees slightly flexed. Lean forward to a 30°–45° angle relative to the ground. Hold the dumbbells with a neutral, closed grip while keeping the wrists neutral. ET006: Open grip.
ET007: Grip width is too narrow or too wide.
ET009: Wrist excessively extended or tilted to the side.
ET013: Excessive elbow adduction or abduction.
ET032: Excessive knee flexion or hyperextension.
ET033: Torso angle too shallow or too steep.
AK53: Keep the upper arms close to the sides of the body while allowing the forearms to hang down. Maintain an approximately 90° angle between the upper and lower arms. ET011: Forearms are not perpendicular to the ground.
ET021: Excessive elbow abduction, moving away from the torso.
ET059: The angle between the upper arm and forearm is too small or too large.
Concentric AK54: Exhale. Keep the upper arms still, elbows close to the torso. Concentrate on the triceps to bend at the elbows, lifting the forearms until they align in a straight line with the upper arms, parallel to the floor. ET014: Inhalation.
ET005: Trunk swaying forward and backward, indicating instability.
ET015: Using a knee drive or leaning the torso backward to assist the movement.
ET020: Elbows move forward or backward, upper arms are not stabilized.
ET021: Excessive elbow abduction, moving away from the torso.
ET047: Uneven elevation of the dumbbells on both sides.
ET048: Insufficient dumbbell elevation, resulting a shortened range of motion.
ET060: Forearms not aligned in a straight line with the upper arms.
Eccentric AK55: Inhale. Use controlled tension in the triceps to lower the dumbbells slowly until the forearms are vertical to the floor. ET017: Exhalation.
ET018: Lowering too quickly.
ET049: Insufficient dumbbell descent, resulting a shortened range of motion.
ET050: Uneven descent of the dumbbells on both sides.
ET059: The angle between the upper arm and forearm is too small or too large.
Table 5: Full list of the FLEX error types’ weight.
Error Code Error Description Weight
ET001 Feet positioned too narrowly or too widely. 6
ET002 Excessive arching or collapsing of the lower back. 9
ET003 Excessive forward flexion or excessive backward curvature of the spine. 9
ET004 Shoulder elevation, lack of scapular retraction, or failure to depress the shoulders. 8
ET005 Trunk swaying forward and backward, indicating instability. 5
ET006 Open grip. 6
ET007 Grip width is too narrow or too wide. 6
ET008 Excessive elbow flexion or hyperextension. 6
ET009 Wrist excessively extended or tilted to the side. 9
ET010 The initial barbell position is set too high or too low. 5
ET011 Forearms are not perpendicular to the ground. 6
ET012 Palms facing inward or downward. 4
ET013 Excessive elbow adduction or abduction. 7
ET014 Inhalation. 2
ET015 Using a knee drive or leaning the torso backward to assist the movement. 8
ET016 Excessive forward head movement, obstructing the barbell’s vertical path. 7
ET017 Exhalation. 2
ET018 Lowering too quickly. 5
ET019 Ending position of the barbell is too high or too low. 7
ET020 Elbows move forward or backward, upper arms are not stabilized. 8
ET021 Excessive elbow abduction, moving away from the torso. 8
ET022 Using forward or upward shoulder movement to assist the action. 8
ET023 Elbows move forward or backward, causing the barbell path to deviate from an arc. 7
ET024 Elbows rising much lower than the shoulder joints. 8
ET025 The barbell does not travel vertically relative to the ground. 9
ET026 Palms facing forward or backward. 4
ET027 Changing the elbow angle, causing the dumbbell path to deviate from an arc. 8
ET028 Ending position of the dumbbell is too high or too low. 7
ET029 Elbows move forward or backward, causing the dumbbell path to deviate from an arc. 8
ET030 Excessive wrist supination over-rotating the forearms. 5
ET031 Initial position of the dumbbell is too far forward or backward. 5
ET032 Excessive knee flexion or hyperextension. 9
ET033 Torso angle too shallow or too steep. 7
ET034 During abduction, the scapula elevates and the neck protrudes forward. 6
ET035 Arms extending diagonally backward, lifting beyond the shoulder joint’s horizontal level. 5
ET036 Feet not firmly planted on the ground. 8
ET037 Head, upper back, or glutes lifting off the bench. 9
ET038 The angle between the upper arm and torso is too large or too small. 8
ET039 Insufficient barbell descent, resulting a shortened range of motion. 7
ET040 Hips rise or legs drive off the floor to assist during the press. 7
ET041 Raising the head or pushing the neck forward to assist during the press. 7
ET042 Shrugging the shoulders or shifting forward during the press. 9
ET043 The barbell sways left or right, or one arm dominates the force output. 9
ET044 Insufficient barbell elevation, resulting a shortened range of motion. 7
ET045 Initial position of the dumbbell is too high or too low. 5
ET046 Dumbbells collide or deviate outward. 5
ET047 Uneven elevation of the dumbbells on both sides. 8
ET048 Insufficient dumbbell elevation, resulting a shortened range of motion. 7
ET049 Insufficient dumbbell descent, resulting a shortened range of motion. 7
ET050 Uneven descent of the dumbbells on both sides. 8
ET051 Initial position of the barbell is too far forward or too far back. 5
ET052 Hips rising faster than barbell, causing asynchronous force distribution between legs and hips. 10
ET053 Excessive forward or backward tilt of the shoulder joint. 9
ET054 Knees bend prematurely, causing the barbell to circle around them. 9
ET055 Excessive forward movement of the knees. 10
ET056 Knees caving inward. 10
ET057 Elbows flex before the scapula retracts. 8
ET058 The hip position is set too high or too low. 6
ET059 The angle between the upper arm and forearm is too small or too large. 6
ET060 Forearms not aligned in a straight line with the upper arms. 6
ET061 Feet placed too narrowly or too widely during the squat. 9
ET062 Barbell deviating from the legs. 8
ET063 The knees lock out, with minimal extension in the hip and knee joints. 10